Hi
a rectangle has the following sides
the widths are x^2 & y respectively
the lengths are 4y+8 & 20x-16 respectively
one side must have an odd integer length.
what is the perimeter.
Thanks

Answer by greenestamps(13200)



The two widths are equal, and the two lengths are equal:
y = x^2
4y+8 = 20x-16 --> 4y = 20x-24 --> y = 5x-6

So

x^2 = 5x-6
x^2-5x+6 = 0
(x-3)(x-2) = 0

x could be either 3 or 2; check to see which (if either) satisfies all the conditions of the problem.

If x = 2 then y = x^2 = 4; the dimensions are 2 and 4. Not right -- the problem says one dimension has to be an odd integer.

If x = 3 then y = x^2 = 9; the dimensions are 3 and 9. This is okay; both dimensions are odd integers.

Now you can find the perimeter to answer the question.